annie and belle are separated by 1km. annie begins walking at a pace of 1m/s toward belle. belle waits 1 minute before beginning his walk of 1.5 m/s toward annie. where will they meet relative to annie's initial position
1
Expert's answer
2020-11-10T06:52:45-0500

First, keep in mind the fact that Annie walked at 1 m/s during 1 minute (or 60 sec). Therefore, this distance is


"d_1=vt_1=1\cdot60=60\text{ m}."

The distance left to Belle is


"d_l=1000-60=940\text{ m}."

Then Belle stars walking and their approach speed is 1+1.5=2.5 m/s. The distance between them is decreasing from 940 m. The time it will take for them:


"t_2=\frac{940}{2.5}=376\text{ s}."

Both Annie and Belle will meet after 376 seconds. The distance passed by Annie is


"d_2=vt_2=1\cdot376=376\text{ m}."

Thus, Annie walked 60+376 m = 436 m.